OpenStudy (anonymous):

Carrie has recorded her first CD and needs 400 for an upcoming release party. CD Clone charges $150 for 100 CD's and $297 for 250 CD's Write an equation for the relationship between the numbers of CD's created and the total cost Book says C=.92d + 67 I'm baffled at how they came up with this answer

OpenStudy (anonymous):

$150 for 100 CD's and $297 for 250 CD's You can realize that the cost will have 2 parts. 1 is independent of the amount of CDs Cloned and the second part is dependent. So the function C will equal to C(x)=ax+b (it depends on x, the number of CDs made) You now can write 2 equations with the 2 data $150 for 100 CD's 150=100a+b $297 for 250 297=250a+b Solve these 2 and luckily you should get the given solution!

OpenStudy (anonymous):

Not sure what you mean as you still have 2 variables in each equation

OpenStudy (anonymous):

150=100a+b 297=250a+b 2 equations with 2 unknowns. 297-150=250a-100a+b-b 147=150a a=0.98